OpenStudy (anonymous):

What is the simplified form of (-5a/3)^-3

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \left(\begin{matrix} \frac{-5a}{3} \\ \end{matrix}\right)^{-3} }\) yes?

OpenStudy (solomonzelman):

when you have a fraction raised to a negative power, you flip it, and raise it to a positive power. \(\LARGE\color{black}{ \left(\begin{matrix} \frac{\color{blue}{C}}{\color{red}{D}} \\ \end{matrix}\right)^{-E} =\left(\begin{matrix} \frac{\color{red}{D}}{\color{blue}{C}} \\ \end{matrix}\right)^{E} }\) Okay? So, \(\LARGE\color{black}{ \left(\begin{matrix} \frac{-5a}{3} \\ \end{matrix}\right)^{-3} ~=~?}\)